OpenStudy (anonymous):

find the equation of the straight line passing inclined at an angle of 135 degree to positive x axis and meeting y axis at (0,-5) 135 degree= -1

OpenStudy (anonymous):

@JakeV8 can you explain?

OpenStudy (ajprincess):

Let the equation of the line be y=mx+c  \(\tan\theta=m\) \(\tan135=m\) Nw to find c plug in the values of m, x and y in y=mx+c Here y=-5, x=0 and m=tan135. Now can u find ur equation?